if ln(a+c)+ln(a-2b+c)=2ln(a-c) find the relation between a,b and c

Grade:11

1 Answers

siddharth gupta
9 years ago
apply the properties of logarithm and obtain the result:
ln(a2-2ab+ac+ac-2bc+c2)=ln(a2+c2-2ac)
on equating arguments of log on both sides and solving you will get the equality:
2ac=ab+bc
divide both sides by abc
finally’2/b=1/a+1/c
i.e a,b,c are in h.p.
 WITH REGARDS
SIDDHARTH GUPTA